OpenStudy (anonymous):

Hi, I always seem to have trouble converting mixed percents like 66 2/3 % to a fraction and a decimal. I know it's pretty basic, but I can't remember the formula and I really need help with it! Thank you!

OpenStudy (anonymous):

66 2/3 is 66 + 2/3. To get a percentage, take the numerator and divide by the denominator. That gives you the percentage. 2/3 = .666666666667 so 66 2/3 = 66.667

OpenStudy (anonymous):

The percent is 66 2/3. I need to know the fraction.

OpenStudy (anonymous):

Okay, so I did some reasearch...According to this website, it says to write the percent over 100, so 66 2/3 over 100. The example problem they have is 62 1/ 2 %, and they convert it to a decimal (66.5) and put that over 100 and multiply it by ten. So I have to change 66 2/3 to a decimal before I can find the fraction but I don't know how.

OpenStudy (anonymous):

You can put 66.67/100 and divide = .6667 The other way is from the first example I gave you, where 66 2/3 = 66.667; To find the decimal percentage, you move the decimal point two spaces to the left = .6667. Hope that helps.

OpenStudy (anonymous):

Okay, I think I got it now. Thank you so much!!!